OpenStudy (anonymous):

Is 3 part of the solution set for the inequality six times w is less than or equal to eighteen? Why or why not?

OpenStudy (anonymous):

Yes, 3 is a part of the solution set for the inequality; \[6w \le 18\] Because the next step is to divide both sides by 6 to get w alone; \[\frac{ 6w }{ 6 }\le\frac{ 18 }{ 6 }\] \[18\div6=3\] So, \[w \le 3\] w is less than or equal to 3